Problems 5 through 9 only pleaseC. Regression Analysis In this section, you will analyze the effects of weather conditions on workers productivity. You coll
Answer #1

5)

6 * log (T_rain) - 2 = 6 log(T_not_rain)

6 log(T_rain) - 6 log(T_not_rain) = 2

6 log ( T_rain / T_not_rain) = 2

log ( T_rain / T_not_rain) = 2/6 = 0.3333

hence

33.333 % more Temperature during rainy days

6)

TS = -2/1.2 = - 1.66666

since n = 500

df = n-k-1 = n-4 = 496

t-critical values are -1.96 and 1.96 for 0.05 level

since TS does not lie in critical region

we fail to reject the null hypothesis

7)

p-value= 2 P( t < -1.6666)

= 0.096225    {approximate by normal distribution}

p-value > 0.05

8)


y^= 12 + 6 * 1.61 - 2 * 1 - 0.05 * 20
= 18.66

9)

t = 1.645 for 90%

(-0.05 - 1.645 * 0.01 , -0.05 + 1.645 * 0.01)

= ( -0.06645 ,-0.03355)

Please rate